OpenStudy (anonymous):

Select the best possible first step to solving the system by first eliminating the y variable. 5x – y = 5 7x – 3y = –1 Multiply the first equation by –3. Multiply the second equation by –3. Multiply the first equation by 3. Multiply the second equation by 3.

OpenStudy (anonymous):

It should be 'A' Multiply the first equation by -3 Because -3(5x – y = 5) 15x+3y=-15 -15x+3y=-15 7x – 3y = –1 ======== -8x =-16